A car moves with an initial velocity of 19 m/s due north. Part A Find the velocity of the car after 5.6 s if its acceleration is 1.6 m/s2 due north Part B Find the velocity of the car after 5.6 s if its acceleration is 1.5 m/s2 due south.
Answer #1

Here,

let east is the positive x direction

north is the positive y direction

initial velocity , u = 19 j m/s

a)

acceleration , a = 1.6 j m/s^2

Using first equation of motion

v = u + a * t

v = 19 + 5.6 * 1.6

v = 28 j m/s

the velocity of the car after 5.6 s is 28 m/s north

b)

acceleration , a = -1.5 j m/s^2

Using first equation of motion

v = u + a * t

v = 19 - 5.6 * 1.5

v = 10.6 j m/s

the velocity of the car after 5.6 s is 10.6 m/s north